OpenStudy (aamey):

What is the most apparent nth term of this sequence (assume that n begins with 1): 1/2, -4/3, 9/4, -16/5

OpenStudy (anonymous):

How do you want to start

OpenStudy (aamey):

I don't know where to start

OpenStudy (anonymous):

Whats the pattern for the numerators

OpenStudy (anonymous):

ignoring the negative for now

OpenStudy (aamey):

1+3=4 4+5=9 9+7=16

OpenStudy (anonymous):

true, but whats 1^2, 2^2, 3^2, 4^2

OpenStudy (aamey):

1,4 ,9,16 so is n squared? is it negative n?

OpenStudy (anonymous):

no? you alternate between + and - right?

OpenStudy (anonymous):

the first part is correct i mean

OpenStudy (anonymous):

try and solve the denom and then we'll look at the negative

OpenStudy (anonymous):

in terms of n

OpenStudy (aamey):

n+1 for the denominator

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so whats (-1) raised to an even power?

OpenStudy (aamey):

-1

OpenStudy (anonymous):

(-1)*(-1) = (-1)^2 = (-1)^even

OpenStudy (anonymous):

= 1 right?

OpenStudy (aamey):

(-1)(-1)(-1) = 1 is that what youre asking

OpenStudy (anonymous):

when you multiply -1 times itself an even number of times it become 1, and if you do it an odd amount of times it stays -1 right?

OpenStudy (aamey):

yes

OpenStudy (anonymous):

so you have +, -, +, -

OpenStudy (anonymous):

if you take (-1)^n what pattern is it

OpenStudy (aamey):

it stays -1

OpenStudy (anonymous):

calculate for n=1, n=2, n=...

OpenStudy (anonymous):

for 1/2 n=1, -4/3 n=2, 9/4 n=3

OpenStudy (aamey):

wait so what would the formula be for the numerator?

OpenStudy (anonymous):

n^2

OpenStudy (anonymous):

look at trying to raise (-1) to something to produce the pattern +, -, +, -

OpenStudy (anonymous):

you start with n=1, so what if you do (-1)^n

OpenStudy (aamey):

-1^1 = -1

OpenStudy (anonymous):

but we know that -1 ^ even is positive, can you add something to n to make the exponent even in the first term?

OpenStudy (anonymous):

\[(-1^{n+1})\frac{ n^2 }{ n+1 }, n \epsilon \mathbb{Z}^+\]